OpenStudy (anonymous):

Given f(x) = 2x − 1 and g(x) = 4x − 3, evaluate the composite function. g[f(x)]

OpenStudy (anonymous):

g[f(x)] = g(2x-1) = 4(2x-1)-3

OpenStudy (anonymous):

Simplify it

OpenStudy (anonymous):

8x -1?

OpenStudy (anonymous):

8x-4-3=8x-7

OpenStudy (kirbykirby):

Not quite, it's 8x-7

OpenStudy (anonymous):

Oh thats right because of the negative 4 my bad thanks!